OpenStudy (anonymous):

a 1.50 kg red ball traveling at 3.0m/s collides head on with a 3.50 kg green marble moving 1.5 m/s. What is the final velocity of each?

OpenStudy (anonymous):

That's one massive marble  Like this? If we're assuming a completely elastic collision, then we know two things about the system. Momentum is conserved and kinetic energy is conserved. Two equations for two unknowns. \[ m_{Ball} = m_1 = 1.5kg \\v_{i \ Ball} = u_i = 3m/s \\ v_{f \ Ball} = u_f = ? \\ \ \\ m_{Marble}=m_2=3.5kg \\ v_{i \ Marble} = v_i = -1.5 m/s \\ v_{f \ Marble} = v_f = ? \] \[\underline{\textrm{Conservation of Momentum}} \\ \ \\ \hspace{60px}p_i = p_f \\ \ \\ m_1u_i + m_2v_i = m_1u_f + m_2v_f \\ \ \\ m_2 v_f = \underbrace{(m_1u_i + m_2v_i)}_{\hbox{We know this value;}}_{\hbox{ call it "A"}} -m_1u_f \\ \hspace{130px}A= (1.5kg)(3m/s)+(3.5kg)(-1.5m/s) \\ \ \\m_2^2v_f^2 = (A-m_1u_f)^2 \\ \] \[ \underline{\textrm{Conservation of Energy}} \\ \ \\ \hspace{60px}K_i = K_f \\ \ \\ \frac{1}{2}m_1u_i^2 + \frac{1}{2}m_2v_i^2 = \frac{1}{2}m_1u_f^2 + \frac{1}{2}m_2v_f^2 \\ \ \\ \ \\ m_1u_i^2 + m_2v_i^2 = m_1u_f^2 + m_2v_f^2 \\ \ \\ \ m_2v_f^2 = \underbrace{m_1u_i^2 + m_2v_i^2}_{\hbox{Call this "B"}} - m_1u_f^2 \\ \ \\ \ \hspace{110px} B = (1.5kg)(3.0m/s)^2+(3.5kg)(-1.5m/s)^2 \\ \ \\ \ m_2(m_2v_f^2)=m_2(B- m_1u_f^2) \\ \ \\ \ \\ \ \\ \ \\ \ m_2(m_2v_f^2) = m_2^2v_f^2 \\ \ \\ m_2(B-m_1u_f^2) = (A-m_1u_f)^2 \] Solve for u_f and then you have the key to finding v_f ! I only named A and B so that the final expression wasn't so gigantic ^_^